Physical Chemistry

A compound forms hexagonal close-packed structure. What is the total number of voids
in 0.5 mol of it? How many of these are tetrahedral voids?

ROHITH
10 Years agoGrade

1 Answer

RAKESH CHINDAM
ApprovedApproved Tutor Answer10 Years ago
Number of close-packed particles = 0.5 × 6.022 × 1023 = 3.011 × 1023
Therefore, number of octahedral voids = 3.011 × 1023
And, number of tetrahedral voids = 2 × 3.011 × 1023 = 6.022 ×1023
Therefore, total number of voids = 3.011 × 1023 + 6.022 × 1023 = 9.033 × 1023